# Fast fitting of phylogenetic mixed effects models

### v1

@@ Abstract
Mixed effects models are among the most commonly used statistical methods for the exploration of multispecies data. In recent years, also Joint Species Distribution Models and Generalized Linear Latent Variale Models have gained in popularity when the goal is to incorporate residual covariation between species that cannot be explained due to measured environmental covariates. Few software implementations of such models exist that can additionally incorporate phylogenetic information, and those that exist tend to utilize Markov chain Monte Carlo methods for estimation, so that model fitting takes a long time. In this article we develop new methods for quickly and flexibly fitting phylogenetic mixed models, potentially incorporating residual covariation between species using latent variables, with the possibility to estimate the strength of phylogenetic structuring in species responses per environmental covariate, and while incorporating correlation between different covariate effects. By combining Variational approximations with a reduced rank matrix normal covariance structure, Nearest Neighbours Gaussian Processes, and parallel computation, phylogenetic mixed models can be fitted much more quickly than the current state-of-the-art. Two simulation studies demonstrate that the proposed combination of approximations is not only fast, but also enjoys high accuracy. Finally, we demonstrate use of the method with a real world dataset of wood-decaying fungi.
@@ Introduction
The exploration of species co-occurrence patterns is at the core of contemporary community ecology. Recent years have seen a push in methods that better facilitate ecologists to understand which species co-occur, and why they co-occur (A. R. Ives and Helmus 2011; Braga et al. 2018). Some species co-occur because they thrive in similar environments, which may be due to similar characteristics in terms of functional traits. Species that share evolutionary history often share such physical characteristics, and consequently hypotheses exist that describe the co-occurrence of closely related species (Pagel 1999).
@@
Data on ecological communities is increasingly analysed with multivariate statistical models. Most notably, Joint Species Distribution Models (JSDMs, Pollock et al. 2014) are now frequently applied for the analysis of binary data of multiple species. Such models represent co-occurrence patterns by incorporating residual covariation that cannot be explained due to the measured environment. Generalized Linear Latent Variable Models (GLLVMs, Warton et al. 2015) are a modeling framework that uses latent variables to represent unmeasured drivers of species co-occurrence patterns. GLLVMs have been presented as a model-based method for ordination (Hui et al. 2015; van der Veen et al. 2023), but are also a computationally efficient method for implementing JSDMs.
@@
GLLVMs tend to be challenging to fit for species that do not have many observations. Such species may be referred to as “rare” by ecologists, but we here generally refer to species that have few observations, such as due to insufficient sampling. In order to ameliorate such issues with few data points, and to incorporate the relatedness of species into the analysis of species co-occurrence patterns, JSDMs have been developed that phylogenetically structure species responses to the environment. In such models, species that are more closely related can be predicted to co-occur (Ovaskainen et al. 2017). This has the benefit of “sharing” information across related species, so that species may more accurately be placed in the environment. In fact, with such models it is possible to predict the environmental preferences of extinct species for which no response observations area available at all, by utilizing phylogenetic information.
@@
Three notable software implementations for such phylogenetic random effects models are the Hmsc (Tikhonov et al. 2024), MCMCglmm (Hadfield 2010), and phyr (A. Ives et al. 2020) R-packages. Unfortunately, these software implementations are either slow because they use Markov chain Monte Carlo for estimation (Hmsc and MCMCglmm) or are limited in their functionality otherwise (e.g., phyr supports few datatypes). Since phylogenetic information and community data of species is becoming more readily available thanks to various openscience initiatives, this motivates development of better, and faster, software implementations of community models that incorporate phylogenetic information.
@@
In this article, we present a new statistical approach for fast fitting of phylogenetic random effects models for community ecological data as part of the gllvm R-package. The proposed modeling framework can alternatively be used for when response data are functional traits, or data on individuals (with a pedigree instead of a phylogeny as in MCMCglmm), but we choose to focus here on community data. The package supports a range of datatypes commonly found in community ecological, such as for counts of individuals (Poisson and negative-binomial, potentially zero-inflated), cover classes (ordinal), percentage cover (beta models potentially with zeros or ones, see Korhonen et al. (2024)), and biomass (Tweedie, see Korhonen et al. (2021) and Niku et al. (2017)). Niku et al. (2021) implemented species-specific random effects as part of the fourth corner latent variable model. Phylogenetic information can facilitates better estimation of species responses to the environment, while latent variables can still be included to account for residual covariation. 1) Phylogenetically structuring the random effects, 2) for which we provide functionality to determine strength of the phylogenetic signal per covariate, and 3) facilitate specification of correlation between random effects using the user-friendly lme4 formula interface (Bates et al. 2015)
@@
Using a combination of methodological developments we can fit the models much more quickly than existing software implementations for Phylogenetically structured random effect models. Variational Approximations (VA, Ormerod and Wand 2010) helps to retrieve a closed form approximation to the likelihood, which we combine with a matrix normal covariance structure for the variational distribution, as well as a reduced-rank approximation to further speed it up. Nearest Neighbour Gaussian processes (NNGPs, Finley et al. 2019) facilitate even quicker computation by forming a sparse approximation to the inverse of the phylogenetic covariance matrix. Finally, fitting the models in parallel via the TMB R-package (Kristensen et al. 2015) further reduces computation time.
@@
With extensive simulations we show that the proposed combination of four approximations still allows us to accurately retrieve the model parameters, and that it tends to be faster than application of the Laplace approximation (e.g., as in the phyr R-package). Using data from Abrego et al. (2022a) we demonstrate use of the model on a real dataset.
@@
In this section we present and explain the phylogenetic random effects model, and in the next section we discuss estimation of the parameters.
Let Y represent a multivariate dataset with rows $i=1\ldots n$ that represent the sites at which an ecological community has been surveyed, and the columns $j=1\ldots m$ represent the species found. $\textbf{X}=\begin{bmatrix}\textbf{x}_{1},\textbf{x}_{2},\ldots\textbf{x}_{k}\end{bmatrix}^{\top}$ is a set of $k=1\ldots p$ covariates with an associated $p\times m$ coefficient matrix $\boldsymbol{\beta}$ for the environment that has been observed at each site, which also includes an intercept column. Covariates can be categorical to represent species-specific intercepts under different conditions, or numerical to represent species-specific responses to the environment, i.e., as slope parameters. As in generalised linear models (GLMs) a link function $\text{g}(\cdot)$ connects the linear predictor $\boldsymbol{\eta}$ to the mean of the conditional distribution assumed for observation $y_{ij}$, so that the model generically is
@@
where X includes an a column of ones for the intercepts. This model can additionally be extended with functional traits as in the fourth corner latent variable model developed in Niku et al. (2021), by hierarchically modeling the species responses:
$\boldsymbol{\beta}=\textbf{1}^{\top}\boldsymbol{\beta}_{x}+\textbf{T}\boldsymbol{B}_{tx}+\boldsymbol{\beta}_{\epsilon},$ (2)
@@
where 1 is a $m$-sized vector of ones, $\boldsymbol{\beta}_{x}$ is a $K$-sized vector of community mean responses to the covariates, T is a $m\times t$ matrix of observed functional traits, $\boldsymbol{B}_{tx}$ is a $t\times p$ matrix of fourth corner coefficients that represent environment-trait interactions, and $\boldsymbol{\beta}_{\epsilon}$ is a matrix of $p\times m$ species-specific random effects for the covariates that model species-specific deviations from the community mean response. The species responses $\boldsymbol{\beta}$ can be thought of as latent traits, which are modeled with observed trait covariates. Excluding the trait covariates simplifies the model to, equivalent to assuming that species responses to the environment are not structured by observed traits.
@@
Let C denote the phylogenetic correlation matrix, and $\boldsymbol{\mu}_{k}$ generically the mean of species environmental responses (with or without traits). To phylogenetically structure the random effects we could assume $\boldsymbol{\beta}_{k}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\sigma_{k}^{2}\textbf{C})$, so that $\sigma_{k}^{2}$ represents the rate of evolution, which determines the variance of the random effect. However, unless there is a phylogenetic signal this model is quite restrictive, so phylogenetic random effects models (e.g., as in the phyr R-package) are usually formulated with a second set of random effects that are independent for all species $\boldsymbol{\beta}_{2k}$ with variance $\sigma_{2k}^{2}$. Since $\boldsymbol{\beta}_{1k}$ and $\boldsymbol{\beta}_{2k}$ are independent, we define $\boldsymbol{\beta}_{k}=\boldsymbol{\beta}_{1k}+\boldsymbol{\beta}_{2k}\sim\mathcal{N}[\textbf{0},\sigma^{2}_{k}\{\textbf{C}\rho_{k}+(1-\rho_{k})\textbf{I}_{m}\}]$ where $\rho_{k}=\frac{\sigma_{1k}^{2}}{\sigma^{2}_{1k}+\sigma_{2k}^{2}}$ is the phylogenetic signal parameter per covariate (equivalent to Pagel’s $\lambda$, see Pagel (1999) and Pearse, Davies, and Wolkovich (2023)), and $\sigma^{2}_{k}=\sigma^{2}_{1}+\sigma^{2}_{2}$ is the variance of the sum of the two terms. Consequently, the variances of the original two terms can be retrieved from this alternative model formulation by nothing that $\sigma^{2}_{1k}=\rho_{k}\sigma_{k}^{2}$ and $\sigma^{2}_{2k}=(1-\rho_{k})\sigma^{2}_{k}$. This formulation has the benefit of only including one set of random effects that need to be integrated out of the likelihood (which is usually computationally intensive for many random effects), but the downside that it requires explicit inversion of the covariance matrix for species at each iteration during model fitting (which tends to be computationally intensive for many species).
@@
Covariances between the random effects of different covariates $\boldsymbol{\Sigma}_{r}$, i.e., for the rows of $\boldsymbol{\beta}_{\epsilon}$ can be introduced by instead defining the distribution of the random effects over all species and covariates simultaneously. Let $\textbf{L}_{k}(\boldsymbol{\Sigma}_{k})$ denote the lower cholesky factor of $\boldsymbol{\Sigma}_{k}$, so that instead we parameterize the covariance over the random effects for all species and covariates as
$\text{vec}(\boldsymbol{\beta}_{\epsilon}^{\top})\sim\mathcal{N}[\textbf{0},\begin{Bmatrix}\textbf{L}(\boldsymbol{\Sigma}_{1})&&&\\ &\textbf{L}(\boldsymbol{\Sigma}_{2})&&\\ &&\ddots&&\\ &&&\textbf{L}(\boldsymbol{\Sigma}_{p})\end{Bmatrix}\biggl{\{}\boldsymbol{\Sigma}_{r}\otimes\textbf{I}_{m}\biggr{\}}\begin{Bmatrix}\textbf{L}(\boldsymbol{\Sigma}_{1})^{\top}&&&\\ &\textbf{L}(\boldsymbol{\Sigma}_{2})^{\top}&&\\ &&\ddots&&\\ &&&\textbf{L}(\boldsymbol{\Sigma}_{p})^{\top}\end{Bmatrix}],$ (3)
@@
where $\boldsymbol{\Sigma}_{mp}=\textbf{C}\rho_{p}+(1-\rho_{p})\textbf{I}_{m}$. If instead we assume $\rho_{k}=\rho$, i.e., that the phylogenetic signal is the same for all covariates, equation (3) reduces to equation (4) of Ovaskainen et al. (2017) or equation (11.1-2) of de Villemereuil and Nakagawa (2014), i.e., $\text{vec}(\boldsymbol{\beta}_{\epsilon})\sim\mathcal{N}(\textbf{0},\boldsymbol{\Sigma}\otimes\boldsymbol{\Sigma}_{r})$. In this formulation, $\boldsymbol{\Sigma}_{r}$ controls correlation between covariate effects of the same species, and together with the phylogenetic signal parameters also the correlation between covariate effects of different species. These off-diagonal blocks in the covariance matrix are by extension formulated as an lu decomposition and take the form $\textbf{L}_{k}(\boldsymbol{\Sigma}_{k})\textbf{L}_{l}(\boldsymbol{\Sigma}_{l})^{\top}\Sigma_{kl}$. The lower triangular matrix of this lu decomposition is $\textbf{L}_{k}(\boldsymbol{\Sigma}_{k})\text{Diag}[\text{diag}\{\textbf{L}_{k}(\boldsymbol{\Sigma}_{k})\}^{-1}$, where $\text{Diag}(\cdot)$ constructs a diagonal matrix from a vector and $\text{diag}(\cdot)$ extracts the diagonal of a matrix. The associated upper triangular matrix is then $\text{Diag}[\text{diag}\{\textbf{L}_{k}(\boldsymbol{\Sigma}_{k})\}]\textbf{L}_{l}(\boldsymbol{\Sigma}_{l})^{\top}\Sigma_{kl}$.
@@
For brevity, let $\text{bdiag}\{\textbf{L}(\boldsymbol{\Sigma}_{k})\}$ represent the blockdiagonal matrix of lower cholesky factors as in equation (3). Consequently, the matrix of species associations at sites $i$ and $l$ due to the phylogenetic random effects model with multiple signal parameters is:
$\boldsymbol{\Sigma}_{spp}=(\textbf{x}_{i}^{\top}\otimes\textbf{I}_{m})\text{bdiag}\{\textbf{L}(\boldsymbol{\Sigma}_{k})\}(\boldsymbol{\Sigma}_{r}\otimes\textbf{I}_{m})\text{bdiag}\{\textbf{L}(\boldsymbol{\Sigma}_{k})^{\top}\}(\textbf{x}_{i}\otimes\textbf{I}_{m})$ (4)
@@
The model in equation (1) can be extended in various ways, for example by incorporating unconstrained latent variables as in Niku et al. (2021), or constrained or informed latent variables (van der Veen et al. 2023), potentially with a quadratic response model (van der Veen et al. 2021), with random site effects to account for pseudoreplication (as demonstrated in the following case study) or spatial autocorrelation, or with additional fixed effects, but we have omitted those here for brevity.
@@
Similarly, in its present form the model only incorporates positive species associations. It is possible to adjust the model to additionally incorporate negative species assocations, but we consider such an implementation as a future avenue for research (see also appendix S1).
@@
Formulating the model in terms of a single set of random effects $\boldsymbol{\beta}_{\epsilon}\sim\mathcal{N}(\textbf{0},\boldsymbol{\Sigma})$, with a vector $\Theta$ of all other parameters in the model (such as dispersion parameters for a Tweedie or negative binomial distribution), we have the following marginal log-likelihood function:
@@
$\mathcal{L}(\boldsymbol{\Theta})=\sum\limits_{j=1}^{m}\log\biggl{[}\int\prod\limits_{i=1}^{n}f\biggl{(}y_{ij}|\boldsymbol{\beta}_{\epsilon}\biggr{)}\mathcal{N}\biggl{\{}\text{vec}\biggl{(}\boldsymbol{\beta}_{\epsilon}\biggr{)};\boldsymbol{\Sigma}\biggr{\}}d\boldsymbol{\beta}_{\epsilon}\biggr{]}.$ (5)
@@
To efficiently retrieve a closed-form approximation to the log-likelihood in equation (5), we use VA (Hui et al. 2017). VA introduces hyperparameters (“variational parameters”) that need to be optimised in order to improve the approximation (tightening the lower bound). With many species and random effects this can significantly slow down model fitting. Other methods for fitting generalized mixed-effects models, such as the Laplace approximation, suffer from similar issues with high dimensions. The Laplace approximation requires evaluating the matrix of second derivatives of the joint likelihood for the random effects and the data, which also significantly slows down for a large number of (correlated) random effects. VA has the benefit of allowing for a second approximation by introducing a more simple structure for the VA covariance matrix, so that the size of the optimisation problem is reduced significantly for quicker model fitting.
@@
Various VA covariance structures are possible to formulate, but some require unrealistic assumptions (such as independence of species or random effects), or have a considerably higher number of variational parameters (e.g., as in an unstructured covariance matrix), so that they require significantly longer computational times. In one of the most reduced scenarios, we assume the variational distribution for the random effects to follow a matrix normal distribution, i.e., $q(\boldsymbol{\beta}_{\epsilon})=\mathcal{MN}(\textbf{a},\textbf{A}_{r},\textbf{A}_{m})$ with $\textbf{A}_{r}$ a $p\times p$ VA covariance matrix and $\textbf{A}_{m}$ a $m\times m$ VA covariance matrix. Some different formulations are provided in the supplementary information (appendix S2) and implemented in the gllvm R-package. For the remainder of this article, we limit ourselves to application of the matrix normal case. We show in the simulations below that the choice of a matrix normal VA distribution still enjoys high accuracy, despite its relative inflexibility.

### v2

# Fast fitting of phylogenetic mixed-effects models

@@ Abstract
Mixed-effects models are among the most commonly used statistical methods for the exploration of multispecies data. In recent years, also Joint Species Distribution Models and Generalized Linear Latent Variale Models have gained in popularity when the goal is to incorporate residual covariation between species that cannot be explained due to measured environmental covariates. Few software implementations of such models exist that can additionally incorporate phylogenetic information, and those that exist tend to utilize Markov chain Monte Carlo methods for estimation, so that model fitting takes a long time. In this article we develop new methods for quickly and flexibly fitting phylogenetic mixed-effects models, potentially incorporating residual covariation between species using latent variables, with the possibility to estimate the strength of phylogenetic structuring in species responses per environmental covariate, and while incorporating correlation between different covariate effects. By combining Variational approximations, a sparse approximation to the phylogenetic precision matrix, and parallel computation, phylogenetic mixed-effects models can be fitted much more quickly than the current state-of-the-art. Two simulation studies demonstrate that the proposed combination of approximations is fast and enjoys high accuracy. We explore sensitivty of the approximation to the ordering of species with a real world dataset of wood-decaying fungi.
Word count: $1.1172\times 10^{4}$
@@ Introduction
The exploration of species co-occurrence patterns is at the core of contemporary community ecology. Recent years have seen a push in methods that better facilitate ecologists to understand which species co-occur, and why they co-occur (Ives and Helmus 2011; Braga et al. 2018). Some species co-occur because they thrive in similar environments, which may be due to similar characteristics in terms of functional traits. Species that share evolutionary history often share such physical characteristics, and consequently hypotheses exist that describe the co-occurrence of closely related species (Pagel 1999).
@@
Data on ecological communities is increasingly analysed with multivariate statistical models. Joint Species Distribution Models (JSDMs, Pollock et al. 2014) are now frequently applied for the analysis of binary data of multiple species. JSDMs are a type of Generalized Linear Mixed-effects Model (GLMM, Brooks et al. 2017; Bates et al. 2015) that incorporates residual covariation to better represent species’ co-occurrence patterns. Generalized Linear Latent Variable Models (GLLVMs, Warton et al. 2015) are a modeling framework that leverages dimension reduction methods to more efficiently implement such multispecies models. GLLVMs have also been presented as a model-based method for ordination (Hui et al. 2015; van der Veen et al. 2023), and have been at the forefront of computational advances for quickly fitting statistical models with random effects (Niku et al. 2019; Korhonen et al. 2021; Kidzinski et al. 2022).
@@
Multivariate statistical models tend to be challenging to fit with species that do not have many observations. Such species may be referred to as “rare” by ecologists, but we here more generally refer to species that have few observations, such as due to insufficient sampling. In order to ameliorate such issues with few data points, and to incorporate the relatedness of species into the analysis of species co-occurrence patterns, JSDMs have been developed that phylogenetically structure species responses to the environment. In such models, species that are more closely related can be predicted to co-occur (Ovaskainen et al. 2017). This has the benefit of “sharing” information across related species, so that species may more accurately be placed in the environment. In fact, with such models it is possible to predict the environmental preferences of extinct species for which no response observations are available at all, by utilizing phylogenetic information.
@@
Three notable software implementations for such phylogenetic mixed-effects models are the Hmsc (Tikhonov et al. 2024), MCMCglmm (Hadfield 2010), and phyr (Ives et al. 2020) R-packages. Unfortunately, these software implementations are either slow because they use Markov chain Monte Carlo for estimation (Hmsc and MCMCglmm) or are limited in their functionality otherwise (e.g., phyr supports few datatypes). Since phylogenetic information and community data of species is becoming more readily available thanks to various openscience initiatives [Hinchliff et al. (2015);], this motivates development of better, and faster, software implementations of community models that incorporate phylogenetic information.
@@
In this article, we present a new statistical approach for fast fitting of phylogenetic mixed-effects models for community ecological data as part of the gllvm R-package (Niku et al. 2023). Phylogenetic information can facilitate better estimation of species responses to the environment, while latent variables can potentially be included to account for residual covariation. The proposed modeling framework can alternatively be used for when response data are functional traits, or data on individuals (with a pedigree instead of a phylogeny as in MCMCglmm), but we choose to focus on community data in this article. The software implementation supports a range of datatypes commonly found in community ecology, such as for counts of individuals (Poisson and negative-binomial, potentially zero-inflated), cover classes (ordinal), percentage cover (beta models potentially with zeros or ones, Korhonen et al. 2024), and biomass (Tweedie models, Korhonen et al. 2021; Niku et al. 2017). Niku et al. (2021) implemented species-specific random effects as part of the fourth corner latent variable model. We extend their work by; 1) phylogenetically structuring the random effects, 2) for which we provide functionality to determine strength of the phylogenetic signal per covariate, and 3) facilitate specification of correlation between the random effects for different covariates using the user-friendly lme4 formula interface (Bates et al. 2015, 2025)
@@
Using a combination of methodological developments we can fit the models much more quickly than existing software implementations for phylogenetic mixed-effects models. Variational Approximations (VA, Ormerod and Wand 2010) helps to retrieve a closed form approximation to the likelihood, which we combine with a matrix normal covariance structure for the variational distribution, as well as a reduced-rank approximation to further speed it up. A sparse approximation to the phylogenetic precision matrix, in the form of a band pattern (Wist and Rue 2006) or with Nearest Neighbour Gaussian processes (NNGPs, Finley et al. 2019) facilitates even quicker computation, i.e., by fitting the models using Gaussian Markov Random fields. Recently, Matsuba et al. (2024) also employed NNGPs to quickly and accurately retrieve the inverse of the phylogenetic covariance matrix, although in a Bayesian context and following for a different model formulation. Finally, we fit the models in parallel via the TMB R-package (Kristensen et al. 2015) to further reduce computation time.
@@
With extensive simulations we show that the proposed combination of four approximations still allows us to accurately retrieve the model parameters, and that it tends to be even faster than fitting the models with the Laplace approximation (e.g., as used in the phyr R-package). The approximation error for the phylogenetic precision matrix is sensitive to row-column permutation (Guinness 2018), an issue that we explore using real data on wood-decaying fungi from Abrego et al. (2022a).
@@
In the following sections we use superscripts to distinguish different quantities, subscripts for indices, bold capitalized characters for matrices, bold lowercase characters for vectors, and ordinary characters to denote scalars. In this section we present and explain the phylogenetic mixed-effects model, and in the next section we discuss estimation of the parameters.
Let Y represent a multivariate dataset with observations $y_{ij}$, where the rows $i=1,\ldots,n$ represent the sites at which an ecological community has been surveyed, and the columns $j=1,\ldots,m$ represent the species found. Let $\textbf{X}=\begin{pmatrix}\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{k}\end{pmatrix}$ denote a set of $k=1,\ldots,p$ covariates, where $\textbf{x}_{i}=\begin{pmatrix}x_{i1},x_{i2},\ldots x_{nk}\end{pmatrix}^{\top}$ with an associated $p\times m$ coefficient matrix B and entries $\beta_{jk}$ for the environment that has been observed at each site, usually including an intercept column. Covariates can be categorical to represent species-specific intercepts under different conditions, or numerical to represent species-specific responses to the environment, i.e., as slope parameters. As in generalised linear models (GLMs) a link function $\text{g}(\cdot)$ connects the model to the conditional mean of the distribution assumed for the responses, so that the model generically is
@@
This model can additionally be extended with functional traits as in the fourth corner latent variable model (Niku et al. 2021), by hierarchically modeling the species responses
$\textbf{B}=\textbf{1}^{\top}\boldsymbol{\beta}^{x}+\textbf{T}\textbf{B}^{tx}+\textbf{B}^{\epsilon},$
@@
where 1 is a $m$-sized vector of ones, $\boldsymbol{\beta}^{x}=\begin{pmatrix}\beta^{x}_{1},\beta_{2},\ldots,\beta^{x}_{p}\end{pmatrix}^{\top}$ is a $p$-sized vector of community mean responses to the covariates, T is a $m\times t$ matrix of species covariates (most commonly functional traits), $\textbf{B}^{tx}$ is a $t\times p$ matrix of interaction coefficients with columns $\boldsymbol{\beta}^{tx}_{k}$, and $\textbf{B}^{\epsilon}$ is a matrix of $p\times m$ species-specific random effects for the covariates that model species’ deviations from the community mean response with entries $\beta^{\epsilon}_{jk}$. The B with coefficients can be thought of as latent traits, which are modeled with species covariates. Excluding the species covariates thus simplifies the model to only include “residual” information.
@@
Let C denote the phylogenetic correlation matrix of dimension $m\times m$, and $\boldsymbol{\mu}_{k}=\boldsymbol{\beta}^{x}_{k}+\textbf{T}\boldsymbol{\beta}^{tx}_{k}$ the community mean response to the $k^{th}$ covariate. To phylogenetically structure the random effects we could assume a model with a single set of random effects $\textbf{b}^{a}_{k}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\gamma_{k}^{2}\textbf{C})$, so that the variance of the random effect $\gamma_{k}^{2}$ represents the rate of evolution. However, in the absence of a phylogenetic signal this model is quite restrictive, so phylogenetic random effects models are usually formulated with a second set of random effects that are independent for all species, $\textbf{b}^{2}_{k}$ with zero mean and variance $\delta_{k}^{2}$. Since $\textbf{b}^{a}_{k}$ and $\textbf{b}^{b}_{k}$ are independent, it follows that $\boldsymbol{\beta}_{k}=\textbf{b}^{a}_{k}+\textbf{b}^{b}_{k}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ where $\boldsymbol{\Sigma}_{k}=\sigma^{2}_{k}\{\textbf{C}\rho_{k}+(1-\rho_{k})\textbf{I}^{m}\}$ and $\rho_{k}=\gamma_{k}^{2}/(\gamma^{2}_{k}+\delta_{k}^{2})$ is the phylogenetic signal parameter per covariate (equivalent to Pagel’s $\lambda$, Pagel 1999; Pearse, Davies, and Wolkovich 2023), while $\sigma^{2}_{k}=\gamma^{2}_{k}+\delta^{2}_{k}$ is the variance of the sum of the two terms. Consequently, the variances of the original two terms can be retrieved from this alternative model formulation by nothing that $\gamma^{2}_{k}=\rho_{k}\sigma_{k}^{2}$ and $\delta^{2}_{k}=(1-\rho_{k})\sigma^{2}_{k}$. Equivalently, $\boldsymbol{\beta}_{k}^{\epsilon}\sim\mathcal{N}(\textbf{0},\boldsymbol{\Sigma}_{k})$. The latter formulation has the benefit of only including one set of random effects that needs to be integrated out of the likelihood (which is usually computationally intensive for many random effects), but the downside that it requires explicit inversion of the covariance matrix for species at each iteration during model fitting (which tends to be computationally intensive for many species).
@@
Covariances between the random effects of different covariates $\boldsymbol{\Sigma}^{r}$, i.e., for the rows of $\textbf{B}^{\epsilon}$ can be introduced by instead defining the distribution of the random effects over all species and covariates simultaneously. Let $\textbf{L}(\boldsymbol{\Sigma}_{k})$ denote the lower cholesky factor of the covariance matrix for the $k^{\text{th}}$ covariate effect, so that instead we parameterize the covariance over the random effects for all species and covariates as
$\text{vec}(\textbf{B}^{\epsilon\top})\sim\mathcal{N}\{\textbf{0},\textbf{L}(\boldsymbol{\Sigma}^{r}\otimes\textbf{I}^{m})\textbf{L}^{\top}\},$ (2)
@@
where L is a block-diagonal matrix with $\textbf{L}(\boldsymbol{\Sigma}_{1}),\dots,\textbf{L}(\boldsymbol{\Sigma}_{p})$ as block matrices. In this formulation, $\boldsymbol{\Sigma}^{r}$ controls correlation between covariate effects of the same species, and together with the phylogenetic signal parameters also the correlation between covariate effects of different species. These off-diagonal blocks in the covariance matrix are by extension formulated as a lower-upper (LU) decomposition and take the form $\textbf{L}(\boldsymbol{\Sigma}_{k})\textbf{L}(\boldsymbol{\Sigma}_{l})^{\top}\Sigma_{kl}$, for covariates $k,l=1\ldots p$. The lower triangular matrix of this LU decomposition is $\textbf{L}(\boldsymbol{\Sigma}_{k})\text{Diag}[\text{diag}\{\textbf{L}(\boldsymbol{\Sigma}_{k})\}^{-1}$, where $\text{Diag}(\cdot)$ constructs a diagonal matrix from a vector and $\text{diag}(\cdot)$ extracts the diagonal of a matrix. The associated upper triangular matrix is then $\text{Diag}[\text{diag}\{\textbf{L}(\boldsymbol{\Sigma}_{k})\}]\textbf{L}(\boldsymbol{\Sigma}_{l})^{\top}\Sigma_{kl}$. If instead we assume $\rho_{k}=\rho$, i.e., that the phylogenetic signal is the same for all covariates, equation (2) reduces to equation (4) of Ovaskainen et al. (2017) or equation (11.1-2) of de Villemereuil and Nakagawa (2014), i.e., $\text{vec}(\textbf{B}^{\epsilon})\sim\mathcal{N}(\textbf{0},\boldsymbol{\Sigma}\otimes\boldsymbol{\Sigma}^{r})$.
@@
Consequently, the matrix of species associations at sites $i$ and $i2$ due to the phylogenetic mixed-effects model with multiple signal parameters is:
$\boldsymbol{\Sigma}_{spp,i,i2}=(\textbf{x}_{i}^{\top}\otimes\textbf{I}^{m})\textbf{L}(\boldsymbol{\Sigma}^{r}\otimes\textbf{I}^{m})\textbf{L}^{\top}(\textbf{x}_{i2}\otimes\textbf{I}^{m}).$
@@
The model in equation (1) can be extended in various ways, for example by incorporating unconstrained latent variables as in Niku et al. (2021), or constrained or informed latent variables (van der Veen et al. 2023), potentially with a quadratic response model (van der Veen et al. 2021), with random site effects to account for pseudoreplication (as included in the following case study) or spatial autocorrelation, or with additional fixed effects, but we have omitted those here for brevity.
@@
Similarly, in its present form the model only incorporates positive species associations. It is technically possible to adjust the model to additionally incorporate negative species assocations, but we consider such an implementation as a potential future avenue for research (see appendix S1 for more information).
@@
For brevity, formulating the model in terms of a single set of random effects with covariance matrix $\boldsymbol{\Sigma}$ given in equation (2). With $f(y_{ij}|\boldsymbol{\beta}^{\epsilon}_{j})$ a response distribution such as a Tweedie or negative binomial distribution, and $\Theta$ a vector incorporating of all other parameters in the model (such as scale parameters), we have the following marginal log-likelihood function:
@@
$\mathcal{L}(\boldsymbol{\Theta})=\sum\limits_{i=1}^{n}\log\biggl{[}\int\prod\limits_{j=1}^{m}f\biggl{\{}y_{ij}|\boldsymbol{\beta}^{\epsilon}_{j}\biggr{\}}\mathcal{N}\biggl{\{}\text{vec}\biggl{(}\textbf{B}^{\epsilon}\biggr{)};\boldsymbol{\Sigma}\biggr{\}}d\textbf{B}^{\epsilon}\biggr{]}.$ (3)
Variational Approximations
@@
To efficiently retrieve a closed-form approximation to the log-likelihood in equation (3), we use Variational Approximations (VA, Hui et al. 2017). VA introduces hyperparameters (“variational parameters”) that need to be optimised in order to improve the approximation (tightening the lower bound). With many species and random effects this can significantly slow down model fitting. Other methods for fitting generalized mixed-effects models, such as the Laplace approximation, suffer from similar issues with high dimensions. The Laplace approximation requires evaluating the matrix of second derivatives of the joint likelihood for the random effects and the data, which also significantly slows down for a large number of (correlated) random effects. VA has the benefit of allowing for a second approximation by introducing a more simple structure for the VA covariance matrix, so that the size of the optimisation problem is reduced significantly for quicker model fitting.
@@
Various VA covariance structures are possible to formulate, but some require stronger assumptions (such as independence of species or random effects), or have a considerably higher number of variational parameters (e.g., as in an unstructured covariance matrix), so that they require significantly longer computational times. In one of the most reduced scenarios, we assume the variational distribution for the random effects to follow a matrix normal distribution, i.e., $q(\textbf{B}^{\epsilon})=\mathcal{MN}(\textbf{a},\textbf{A}^{r},\textbf{A}^{m})$ with a a $p\times m$ matrix of variational mean parameters, $\textbf{A}^{r}$ a $p\times p$ VA covariance matrix, and $\textbf{A}^{m}$ a $m\times m$ VA covariance matrix. Some different formulations are provided in the supplementary information (appendix S2) and implemented in the gllvm R-package. For the remainder of this article, we limit ourselves to application of the matrix normal case, with unstructured or diagonal $\textbf{A}^{r}$. We show in the simulations below that the choice of a matrix normal VA distribution still enjoys high accuracy, despite its relative inflexibility, even with diagonal $\textbf{A}^{r}$.